Assume the frictional force on the boat is equal to 1804 N, and the boat is accelerating at a rate of 1.20 m/s2. What force does the truck need to pull on the boat- counteracting the friction- to accelerate it if the trailer (boat) has a mass of 300 kg
1
Expert's answer
2020-12-17T09:03:41-0500

According to the second Newtons low


"F-F_f=ma\to F=ma+F_f=300\cdot1.2+1804=2164(N)"